The sum of 2 numbers is 20 and their product is 75. What are the 2 numbers?
Found 2 solutions by jorel555, stanbon:
Answer by jorel555(1290)

Let the two numbers be m and n. Then:
m+n=20
n=20-m
m(20-m)=75
mē-20m+75=0
(m-15)(m-5)=0
m=15,5
Your two numbers are 15 and 5. ☺☺☺☺

Answer by stanbon(75887)

The sum of 2 numbers is 20 and their product is 75. What are the 2 numbers?
-----
x + y = 20
x*y = 75
----------
(20-y)y = 75
20y - y^2 = 75
------
y^2 - 20y + 75 = 0
(y-15)(y-5) = 0
----
If y = 15, x = 5
-----------------
Cheers,
Stan H.
-----------